OpenStudy (anonymous):

The geometric mean of x and 5 is 15. Find the value of x.

OpenStudy (anonymous):

\[x \times 5= 15 \]

OpenStudy (anonymous):

\[x + 5 = 15?\]

OpenStudy (anonymous):

x+5=15 is 10 x times 5 =15 is 3

OpenStudy (anonymous):

x = 3 because 15 divided by 5 is 3?